a line passes through a point with position vector i(cap)-2j(cap)-k(cap) and is parallel to the vector i(cap)-2j(cap)+2k(cap) find the distance of point p(5,0,-4) from the line and how?

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
The equation of line with the given conditions :
(x - 1)/1 = (y+2)/-2 = (z+1)/2
Now calculate the distance of this line from the point p by distance formula. Simple
Thanks
Bharat Bajaj
askiitians Faculty
IIT Delhi